Consider the complex number =e^(i π/3) . Write that complex number in rectangular form
Answer by jim_thompson5910(35256)

e^(i π/3) = cis(π/3)

e^(i π/3) = cos(π/3) + i*sin(π/3)

e^(i π/3) = 1/2 + i*(sqrt(3)/2)

e^(i π/3) = 1/2 + (sqrt(3)/2)*i

Answer: e^(i π/3) = 1/2 + (sqrt(3)/2)*i